# Denoising Vision Transformers

### v1

@@ 4.3 Generalizable Denoiser
Our per-image denoising method can already effectively remove artifacts from ViT outputs, yielding visually stunning denoised feature maps, as showcased in Figure 1. The problems we are left with are run-time efficiency and distribution shifts. Specifically, the per-image approach is suboptimal for real-time applications, and individually denoised feature maps can lead to distribution shifts due to sample bias, which hampers the feature coherence across different images. To address these issues, we introduce a generalizable denoiser.
After per-image denoising, we accumulate a dataset of pairs of noisy ViT outputs $\mathbf{y}$ and their denoised counterparts $\mathcal{F}$, denoted as =$\mathcal{B}=\{\left(\mathbf{y}_{i},\mathcal{F}_{i}\right)\}|_{i=1}^{B}$. To achieve a generalizable denoising model, we distill these individually denoised samples into a denoiser network ${D}_{\zeta}$, which is trained to predict noise-free features from raw ViT outputs. The training objective is formulated as:
@@
Specifically, our generalizable denoiser consists of a single Transformer block, supplemented with additional learnable positional embeddings that are applied post the forward pass of a ViT. This design aims to mitigate the input-independent artifacts. To predict denoised features, the outputs from a pre-trained ViT are added with these positional embeddings and then processed through the Transformer block. This can be efficiently implemented in a single line of code:
Here, self.denoiser refers to the single Transformer block, and self.PE represents the additional learnable positional embeddings, and y is the ViT output. Notably, this learned denoiser is lightweight, thus adding minimal latency to the original ViT. It also learns to generalize across samples, enabling real-time applications and mitigating the distribution shift issue inherent to per-image denoising.
@@ 5 Experiments
In this section, we first test our per-image denoising algorithm on ViTs trained with different objectives. Then, we evaluate the effectiveness of our generalizable denoiser on dense prediction tasks. For all experiments, we default to using ViT-base models with patch sizes of 14 or 16, depending on the availability of their implementations and model weights in PyTorch Image Models (timm [37]). We defer the implementation details to the supplementary material.
@@ 5.1 Artifacts in ViTs
First, we explore if ViTs trained with different objectives exhibit similar artifacts. To this end, we test with a few representative ViTs, categorizing them into two groups based on the severity of observed artifacts: one with strong artifacts and the other with mild artifacts.
Algorithms producing strong artifacts.
We highlight several ViT training algorithms that result in pronounced feature artifacts, as observed in Figure 1 (except for (d)). Among these, DINOv2 [22], a state-of-the-art vision foundation model with excellent performance on downstream tasks, displays clear position-related artifacts. Additionally, DeIT-III [32], trained with class labels, and CLIP [26], trained by text-image alignment, also exhibit noticeable artifacts. Furthermore, EVA02 [13], which distills local patch features from a pre-trained CLIP model using masked image modeling, also has clear feature artifacts. Our proposed method successfully mitigates these artifacts in all the tested ViTs (compare “Original” and “Denoised” in Figure 1).
Algorithms producing mild artifacts. Conversely, certain models demonstrate only weak artifacts. Specifically, DINO [4] and MAE [15] tend to exhibit low-frequency patterns that are less visually noticeable in individual images [^2] These patterns are more prominent in videos.. Intriguingly, while DINOv2 [22] trained with register tokens (DINOv2-reg [8]) initially appears to be free from artifacts in [8], our DVT uncovers their existence (Figure 1-(d), and its bottom row). Although DINOv2-reg shows fewer artifacts compared to the standard DINOv2, it still displays more artifacts than DINO and MAE. We recognize Register as an improved ViT training technique, but it does not fundamentally eliminate the artifacts.
Table 1: Comparison of features correlation to spatial positions. We report the maximal information coefficient (MIC) between grid features and their normalized patch coordinates.
@@
Original Artifacts Semantics
@@
Correlation between artifacts and positions. Beyond qualitative analyses, we quantitatively investigate the correlation between artifacts and patch positions. Specifically, we compute the maximal information coefficient (MIC) between grid features and their normalized patch coordinates (elaborated in the Appendix). This metric indicates the correlation extent between features and spatial positions. Table 1 presents the results. We observe that both the original ViT outputs and the decomposed artifacts exhibit a stronger spatial correlation than the denoised semantic features, regardless of the training approach. This confirms the link between positional embeddings and the emergence of undesirable artifacts.
Table 2: Qualitative performance of DVT. DVT improves differently pre-trained ViTs for dense prediction tasks. We report performance on semantic segmentation (VOC2012, ADE20K) and depth prediction (NYUd) tasks. The best results are bolded.
VOC2012 [11] ADE20k [40] NYUd [21]
mIoU ($\uparrow$) aAcc ($\uparrow$) mAcc ($\uparrow$) mIoU ($\uparrow$) aAcc ($\uparrow$) mAcc ($\uparrow$) RMSE ($\downarrow$) Rel ($\downarrow$)
Weak Artifacts
MAE [15] 50.24 88.02 63.15 23.60 68.54 31.49 0.6695 0.2334
MAE [15] + DVT 50.53 88.06 63.29 23.62 68.58 31.25 0.7080 0.2560
DINO [4] 63.00 91.38 76.35 31.03 73.56 40.33 0.5832 0.1701
DINO [4] + DVT 66.22 92.41 78.14 32.40 74.53 42.01 0.5780 0.1731
DINOv2-reg [8] 83.64 96.31 90.67 48.22 81.11 60.52 0.3959 0.1190
DINOv2-reg [8] + DVT 84.50 96.56 91.45 49.34 81.94 61.70 0.3880 0.1157
Strong Artifacts
DeiT-III [32] 70.62 92.69 81.23 32.73 72.61 42.81 0.5880 0.1788
DeiT-III [32] + DVT 73.36 93.34 83.74 36.57 74.44 49.01 0.5891 0.1802
EVA02 [13] 71.52 92.76 82.95 37.45 72.78 49.74 0.6446 0.1989
EVA02 [13] + DVT 73.15 93.43 83.55 37.87 75.02 49.81 0.6243 0.1964
CLIP [26] 77.78 94.74 86.57 40.51 76.44 52.47 0.5598 0.1679
CLIP [26] + DVT 79.01 95.13 87.48 41.10 77.41 53.07 0.5591 0.1667
DINOv2 [22] (reprod.) 83.60 96.30 90.82 47.29 80.84 59.18 0.4034 0.1238
DINOv2 [22] + DVT 84.84 96.67 91.70 48.66 81.89 60.24 0.3943 0.1200
@@ 5.2 Evaluation on Downstream Task Performance
Setup. We follow [22, 8] to assess our denoiser across several benchmarks: semantic segmentation tasks on VOC2012 [11] and ADE20k [40], and the depth prediction task on the NYU-depth benchmark [21], using a linear probing protocol. It is important to note that there is no direct competitor for these tasks in our study. Instead, our focus is on comparing the performance of pre-trained ViTs before and after applying our DVT. For all the models in the main experiments, we use 10k denoised samples randomly selected from the VOC2012 and the VOC2007 datasets, excluding their validation samples, to train the second-stage denoiser.
@@
Results. Table 2 presents the main results. We observe significant and consistent enhancements in nearly all pre-trained ViTs across various dense prediction tasks post-denoising. These improvements are achieved without expensive re-training of ViTs at scale, unlike Register [8]; our DVT uses just a single Transformer block for denoising. Notably, the DINOv2-giant model, with an 83.0 mIoU on VOC2012 as reported in [22], is significantly outperformed by our DVT-denoised DINOv2-base model (84.84 mIoU). This improvement extends to the ADE20k dataset, where the DINOv2-giant and DINOv2-large models yield mIoUs of 49.0 and 47.7, respectively as in [22], while our denoised base model achieves a 48.66 mIoU. These results suggest that the performance enhancement is primarily due to effective artifact removal, rather than the tiny increase in the number of parameters of our denoiser network.
@@
Enhancement of DINOv2 with register tokens. Our DVT also boosts the performance of the recently introduced DINOv2-reg model [8], where a ViT is trained with dummy learnable register tokens. As shown in Table 2, our DVT significantly enhances the performance of both DINOv2 [22] and DINOv2-reg [8]. When applying DVT only, DINOv2 witnesses more improvements compared to using registers; for instance, DINOv2 denoised by DVT achieves 84.84 mIoU in VOC2012 and 48.66 mIoU in ADE20k, surpassing the performance of DINOv2-reg, which attains 83.64 mIoU and 48.22 mIoU on the respective benchmarks. Additionally, DVT can further enhance the performance of DINOv2-reg [8] by a substantial margin on both datasets (+0.86 in VOC2012 and +1.12 in ADE20k). These findings suggest that DVT is more adept at addressing the artifact issue inherent in ViTs. In addition, DINOv2-reg [8] requires training ViTs from scratch using 142M images, while our approach only requires training a single Transformer block using 10k denoised samples.
@@
Figure 5: Emerged object discovery ability. We present qualitative results for DVT’s learned denoiser outputs. Features are visualized using PCA and $L2$ feature norms, comparing original ViT features with our denoised features across different algorithms. Noticeably, DVT denoised features show higher feature norms on objects of interest and reduced high- (see a, b) and low-norm artifacts (see c, d).
5.3 Qualitative results
Visual analysis of ViTs.
In Figure 4, we present a visual analysis of the artifact decomposition across various layers of DINOv2 ViTs of different sizes (b), alongside feature maps generated using only zero-tensors as input (a). Notably, the artifacts decomposed by our DVT show a strong visual resemblance to these zero-tensor-input feature maps. In addition, we observe that the artifacts vary across layers: the shallower layers predominantly exhibit low-frequency patterns, whereas the deeper layers are characterized by high-frequency patterns. Importantly, these patterns are consistent across ViTs of different sizes (e.g., from ViT-small to ViT-large), contradicting the suggestion in [8] that only large and sufficiently trained ViTs would display such patterns. Further, Figure 4-(c) showcases the enhanced similarity of central patches compared to other patches post-denoising. Lastly, we see that the artifacts in feature maps will hurt the K-means clustering accuracy significantly and our DVT addresses this issue. These factors are particularly important for dense prediction tasks.
Emerged object discovery ability.
@@ 5.3 Qualitative results
An intriguing finding from our experiments is the emerging capability of object discovery in denoised ViTs. Figure 5 illustrates this through PCA visualizations and $L2$ norms of the feature maps. Post-denoising, not only are the artifacts removed, but also the objects of interest become more distinctly visible. This enhancement in object clarity is not an original goal of DVT but emerges as the outcome of our method. It is noteworthy that not all pre-trained ViTs initially demonstrate this object discovery ability, as seen in Figure 5-(b,c,d) “Original PCA”; however, this capability is remarkably evident after the denoising process. It intriguingly implies an intrinsic property of denoised ViTs — finding salient objects.
5.4 Ablation Study
In this section, we provide ablation studies to understand the importance of different components in our proposed DVT. We use DINOv2-base [22] for the experiments here.
@@ 5.4 Ablation Study
Table 3: Ablation study on per-image denoising using KNN segmentation evaluation protocol on the VOC2012 validation set.
@@
Factorization.
We ablate our per-image denoising method using a K-Nearest-Neighbor (KNN) pixel segmentation evaluation protocol on the VOC2012 dataset. Specifically, we collect class centroids from each training image by masked pooling to construct a memory bank using ground truth annotations. Then, for each pixel in a validation image, we classify it based on its 20 nearest neighbors in the memory bank. We report the mIoU on the validation set. Table 3 shows the results. We observe that combining the artifact field $\mathcal{G}$ and the residual term $\hat{\Delta}$ yields the best result (d). Omitting both these elements reduces our approach to merely utilizing a neural field $\mathcal{F}$ to learn multi-crop ensembled image features, without addressing artifacts (b). While this variant shows improvement, it falls behind our proposed method by a large margin, underscoring the importance of removing artifacts.
@@
We explore alternative architectural designs for our generalizable denoiser in Table 4. We study four variations: 1) our default setting, which incorporates a single Transformer Block with new learnable position embeddings; 2) our default setting but without position embeddings; 3) a multi-layer convolution denoiser with a Conv1x1-ReLu-Conv1x1-ReLu-Conv1x1 structure, and 4) a multi-layer convolution denoiser with a Conv3x3-ReLu-Conv3x3-ReLu-Conv3x3 structure. We observe that the denoisers based on convolutional structures (b, c) do not yield good results, with the conv1x1 setting performing the worst (c). Moreover, we note that our default setting with a Transformer block and learnable positional embedding achieves the best result (d), and removing learnable position embeddings obtains similar numerical performance (e), but we find that our default setting (Transformer Bloack + PE.) is more sensitive to local details such as text and watermark, as shown in Figure S7. Additionally, qualitative comparisons in Figure S7 highlight that convolution-based denoisers typically struggle with removing artifacts.

### v2

@@ 4.3 Generalizable Denoiser
Our per-image denoising method can already effectively remove artifacts from ViT outputs, yielding visually stunning denoised feature maps. The problems we are left with are run-time efficiency and distribution shifts. Specifically, the per-image denoising process is suitable for offline applications but undesired for real-time applications, and individually denoised feature maps can lead to feature distribution shifts due to sample bias, which hampers the feature coherence across images. To address these issues, we introduce a generalizable denoiser.
After applying per-image denoising, we accumulate a dataset of pairs consisting of noisy ViT outputs $\mathbf{y}$ and their denoised counterparts $\mathcal{F}$, denoted as =$\mathcal{B}=\{\left(\mathbf{y}_{i},\mathcal{F}_{i}\right)\}_{i=1}^{B}$. We then train a denoiser network ${D}_{\zeta}$ to predict noise-free features from raw ViT outputs, i.e., ^$\hat{\mathcal{F}}=D_{\zeta}(\mathbf{y})$. The loss function is:
@@
Our generalizable denoiser is implemented as a single Transformer block, supplemented with additional learnable positional embeddings that are applied post the forward pass of a ViT. This design aims to mitigate the input-independent artifacts. To predict denoised features, the outputs from a pre-trained ViT are added with these positional embeddings and then processed through the Transformer block.
Notably, this learned denoiser is lightweight, thus adding negligible latency to the original ViT and facilitating real-time applications. It also learns to generalize across samples, mitigating the distribution shift issue in the per-image denoising process.
@@ 5 Experiments
In this section, we first explore if ViTs trained with different objectives all have artifacts. Then, we evaluate the effectiveness of our generalizable denoiser on dense prediction tasks. For all experiments, we default to using ViT-base models with patch sizes of 14 or 16, depending on the availability of their implementations and model weights in PyTorch Image Models (timm [42]). We defer all the implementation details to the appendix.
@@ 5.1 Artifacts in ViTs
Positional artifacts in different ViTs.
We visualize feature maps from differently pre-trained ViTs in Fig. 1. Among these, DINOv2 [25], a state-of-the-art vision foundation model with excellent performance on downstream tasks, displays clear position-related artifacts. Additionally, DeIT-III [36], trained with image class labels, and CLIP [27], trained by text-image alignment, also exhibit noticeable artifacts. Furthermore, EVA02 [13], which distills local patch features from a pre-trained CLIP model using masked image modeling, also has clear feature artifacts. In ViTs we have tested, our proposed DVT successfully mitigates these artifacts (“Original features” vs. “Denoised features” in Fig. 1).
Figure 5: Visual analysis of ViT output features and denoised features. (a) Visualizations of the feature maps from all layers of a DINOv2 [25] ViT-base model. Notably, the artifacts in the feature maps derived from the cat image exhibit a strong visual correlation with those from the zero-tensor inputs. (b) Visualizations of the decomposed artifacts, the original features, and the denoised features across various layers of DINOv2 ViTs. We observe similar patterns in differently-sized models.
Artifacts in different layers.
In Fig. 5, we present a visual analysis of the artifact decomposition across various layers of DINOv2 ViTs of different sizes (b), alongside feature maps generated using only zero-tensors as input (a). Notably, the artifacts decomposed by our DVT show a strong visual resemblance to these zero-tensor-input feature maps. In addition, we observe that the artifacts vary across layers: the shallower layers predominantly exhibit low-frequency patterns, whereas the deeper layers are characterized by high-frequency patterns. Importantly, these patterns are consistent across ViTs of different sizes (e.g., from ViT-small to ViT-large), diverging from the hypothesis in [7] that only large ViTs would display such patterns.
Table 1: Comparison of features correlation to spatial positions. We report the maximal information coefficient (MIC) between grid features and their coordinates.
@@
Method Raw features Artifacts term $\mathcal{G}$ Semantics term $\mathcal{F}$
@@
Correlation between artifacts and positions.
Beyond visual qualitative inspection, we aim to quantitatively analyze the correlation between artifacts and their positions. Similar to [40], we use the maximal information coefficient (MIC) to measure the dependency between grid features and their normalized patch coordinates (See appendix for more details). This metric indicates how much patch features depend on their spatial positions and semantic content. As shown in Tab. 1, both the original ViT outputs and the decomposed artifacts exhibit a higher spatial correlation than the denoised semantic features, irrespective of the training methodology employed. These results support our hypothesis about the significant role of positional embeddings in the emergence of artifacts. Note that there is no “ground-truth” quantitative metric to to definitively quantify these patterns; hence, our reported numerical results should be viewed as empirical indicators, akin to the “high-norm” indicator used in [7].
@@ 5.2 Evaluation on Downstream Task Performance
We evaluate our method in dense recognition tasks, including semantic segmentation, monocular depth estimation, object detection, and object discovery. It is important to note that there is no direct competitor for these tasks in our study. Instead, our focus is on comparing the performance of pre-trained ViTs before and after applying our DVT. For all the models in the main experiments, we use 10k denoised samples randomly selected from the VOC2012 and the VOC2007 datasets, excluding their validation samples, to train generalizable denoisers.
Semantic segmentation.
@@
We follow [25, 7] to evaluate our approach in two semantic segmentation datasets: VOC2012 [12] and ADE20k [45], using a linear probing protocol, i.e., a linear layer is trained to predict pixels’ class from patch tokens. Tab. 2 presents the main results. We observe significant and consistent enhancements in all pre-trained ViTs across datasets. Notably, the DINOv2-giant, with an 83.0 mIoU on VOC2012 as reported in [25], is outperformed by our DVT-denoised DINOv2-base model (84.84 mIoU). This improvement is also evident in the ADE20k dataset, where the DINOv2-giant and DINOv2-large models attain mIoUs of 49.0 and 47.7, respectively, as reported in [25], while our denoised base model achieves a 48.66 mIoU. Remarkably, the giant model, which is $\mathbf{13\times}$ larger than the base model, is outperformed by or on par with our denoised base model. This indicates that the performance gains primarily stem from effective artifact removal rather than the minor increase in model parameters of our denoiser network.
@@
Our DVT also increases the performance of the concurrent DINOv2-reg model [7], where a ViT is trained with dummy learnable register tokens. As evidenced in Tab. 2, our DVT enhances the performance of both DINOv2 ((f1) vs. (f2)) and DINOv2-reg ((e1) vs. (e2)). When applying DVT only, DINOv2 shows more improvements compared to using registers ((f2) vs. (e1)); for instance, DINOv2 denoised by DVT achieves 84.84 mIoU in VOC2012 and 48.66 mIoU in ADE20k, surpassing the performance of DINOv2-reg, which achieves 83.64 mIoU and 48.22 mIoU on the respective benchmarks. Furthermore, DVT can further enhance the performance of DINOv2-reg ((e1) vs. (e2)) on both datasets (+0.86 in VOC2012 and +1.12 in ADE20k). In addition, DINOv2-reg [7] requires retraining entire models from scratch using 142M images, while our approach requires training a single Transformer block using 10k denoised samples.
Table 2: Quantitative performance of DVT. DVT improves differently pre-trained ViTs for dense prediction tasks. We report performance on semantic segmentation (VOC2012, ADE20K) and depth prediction (NYUd) tasks.
VOC2012 Segmentation ADE20k Segmentation NYUv2 Depth Estimation
Method mIoU($\uparrow$) mAcc($\uparrow$) mIoU($\uparrow$) mAcc($\uparrow$) $\delta_{1}$($\uparrow$) abs rel($\downarrow$)
(a1) DINO [3] 63.00 76.35 31.03 40.33 73.19 0.1701
(a2) DINO [3] + DVT 66.22 78.14 32.40 42.01 73.53 0.1731
(b1) DeiT-III [36] 70.62 81.23 32.73 42.81 72.16 0.1788
(b2) DeiT-III [36] + DVT 73.36 83.74 36.57 49.01 71.36 0.1802
(c1) EVA02 [13] 71.52 82.95 37.45 49.74 63.68 0.1989
(c2) EVA02 [13] + DVT 73.15 83.55 37.87 49.81 68.52 0.1964
(d1) CLIP [27] 77.78 86.57 40.51 52.47 73.95 0.1679
(d2) CLIP [27] + DVT 79.01 87.48 41.10 53.07 74.61 0.1667
(e1) DINOv2-reg [7] 83.64 90.67 48.22 60.52 87.88 0.1190
(e2) DINOv2-reg [7] + DVT 84.50 91.45 49.34 61.70 88.26 0.1157
(f1) DINOv2 [25] 83.60 90.82 47.29 59.18 86.88 0.1238
(f2) DINOv2 [25] + DVT 84.84 91.70 48.66 60.24 87.58 0.1200
Depth estimation.
Following [25], we evaluate our method on the NYUv2-Depth dataset [24] using a linear evaluation protocol (more details in appendix). As shown in Tab. 2, our method clearly enhances the performance of most pre-trained ViTs. For context, the DINOv2-large model exhibits a 0.01 RMSE improvement over the DINOv2-base model with 3.5$\times$ more parameters. Our denoiser achieves similar performance gains with 0.08$\times$ the parameters of the base model. These results highlight our method’s efficiency, achieving marked performance gains with minimal increases in parameter count.
Table 3: Object detection with frozen features. We report the mAP metric on the VOC object detection benchmark.
Method DINOv2 [25] DINOv2-reg [7] CLIP [27] DeiT-III [36] DINO [3] EVA02 [14]
baseline 81.4 80.9 80.9 75.8 76.4 79.4
baseline + DVT 81.9 (+0.5) 81.4 (+0.5) 81.7 (+0.9) 77.0 (+1.2) 77.1 (+0.7) 80.2 (+0.8)
Object detection.
In this experiment, we train ViTDet detectors [21] on the frozen features following the Faster RCNN framework [31] (more details in appendix). We train all models on the VOC trainval07+12 subset and report their mAP metrics on the test2007 subset. Results are reported in Tab. 3. Our approach shows consistent improvements over the studied ViTs. Notably, DINOv2-reg [7] shows a slight decrease in object detection performance when compared to the original DINOv2 [25], while our approach improves it.
@@
Figure 6: Emerged object discovery ability. The features denoised by our DVT show higher feature norms on objects of interest.
Table 4: Unsupervised object discovery using LOST [33]. We report the corloc score across three datasets. Our DVT significantly improves existing models. : results quoted from [7]; these models are ViT-large trained on the ImageNet-22k dataset while our reported results are based on the publicly available ViT-base.
Method VOC2007 VOC2012 COCO20k
(a) DINOv2 [25] 35.3 40.2 26.9
(b) DINOv2-reg [7] 55.4 60.0 42.0
(c) DINOv2-reg 38.0 41.5 26.9
(d) DINOv2-reg + DVT 56.1 (+18.1) 59.3 (+17.8) 45.5 (+18.6)
(e) DINOv2 30.8 35.9 23.4
(f) DINOv2 + DVT 58.0 (+27.2) 60.3 (+24.4) 46.7 (+23.3)
Object discovery.
@@
Unsupervised object discovery has been a long-standing problem of interest. An intriguing finding from our experiments is the emerging capability of object discovery in denoised ViTs. Fig. 6 illustrates this through PCA visualizations and $L2$ norms of the feature maps. Post-denoising, not only are the artifacts removed, but also the objects of interest become more distinctly visible from the feature norm values. This enhancement in object clarity is not a goal of DVT but emerges as the outcome of our method.
To quantitatively assess these enhancements, we follow [7] to use LOST [33] for evaluating object discovery efficacy before and after applying our DVT. We use feature norms as an indicator of object prominence. We conduct object discovery experiments on PASCAL VOC 2007 [11] and 2012 [12] and COCO20k datasets [22]. Tab. 4 presents the results. Our DVT significantly improves both DINOv2 [25] and DINOv2-reg [7] in all the evaluated datasets. In particular, while the publicly available DINOv2-reg shows some improvements ((c) vs. (e)), we find that it falls short of the performance levels reported in [7] ((c) vs. (b)). Despite this, our DVT achieves more substantial enhancements in object discovery capabilities, even compared to the numbers reported in [7] ((f) vs. (b)).
@@
Table 5: Ablation study on per-image denoising using KNN segmentation protocol on VOC12val.
@@
5.3 Ablation Study
In this section, we provide ablation studies to understand the importance of different components in our proposed DVT.
Factorization.
We ablate our per-image denoising method using a K-Nearest-Neighbor (KNN) pixel segmentation evaluation protocol on the VOC2012 dataset. Specifically, we collect class centroids from each training image by masked pooling to construct a memory bank using ground truth annotations. Then, for each pixel in a validation image, we classify it based on its 20 nearest neighbors in the memory bank. We report the mIoU on the validation set. Sec. 5.2 shows the results. We observe that combining the artifact field $\mathcal{G}$ and the residual term $\hat{\Delta}$ yields the best result (d). Omitting both these elements reduces our approach to merely utilizing a neural field $\mathcal{F}$ to learn multi-crop ensembled image features, without addressing artifacts (b). While this variant shows improvement, it falls behind our proposed method by a large margin, underscoring the importance of removing artifacts.
@@ 5.3 Ablation Study
We explore alternative architectural designs for our generalizable denoiser in Sec. 5.2. We study four variations: 1) our default setting, which incorporates a single Transformer Block with new learnable position embeddings; 2) our default setting but without position embeddings; 3) a multi-layer convolution denoiser with a Conv1x1-ReLu-Conv1x1-ReLu-Conv1x1 structure, and 4) a multilayer convolution denoiser with a Conv3x3-ReLu-Conv3x3-ReLu-Conv3x3 structure. We observe that denoisers based on convolutional structures (b, c) do not yield good results, with the conv1x1 setting performing the worst (c). Moreover, we note that our default setting with a Transformer block and learnable positional embeddings achieves the best result (d), and removing the learnable position embeddings obtains very similar numerical performance (e). We empirically find that the design of (d) leads to better qualitative visualizations, and thus we use this setting.
Figure 7: DVT’s Scaling Behaviors. We study the generalizable denoiser’s performance for (a) different model sizes, (b) the number of denoised samples used for training denoisers, and (c) the number of views used when performing per-image denoising.
Scaling behaviors.
We study how DVT scales with model sizes and data scales in Fig. 7. In (a), we see DVT boosts differently-sized ViTs, even allowing ViT-base to match or exceed ViT-giant performance in semantic segmentation. Overall, DVT’s scaling behaviors closely align with those of baseline models. In (b), we study the impact of the number of denoised training samples on task performance, where DVT shows promising results even with limited training samples (e.g., 100$\sim$1000). Note that our denoiser never sees ADE20k and NYU-depth datasets during training, yet generalizes effectively. In (c), we plot the task performance vs. the number of views used for the denoising. DVT benefits from more views in first-stage denoising. When training neural fields, more views enhance performance, while fewer views lead to overfitting. In particular, aggregating views is itself an approach to denoising, which still aligns with our motivation. We also demonstrate that a denoiser trained on samples denoised solely by aggregating views via neural fields (${\mathcal{F}}$-only in (c)) surpasses baselines but underperforms the full DVT, which further confirms the effectiveness of our proposed denoising procedure.